OpenStudy (anonymous):

How do i factor this ? Using complete sentences, explain how to completely factor 3x2 − 3x − 60.

pooja195 (pooja195):

to factor 3x2 − 3x − 60 you have to divide the whole equation by 3 because x^2 cannot be 3x^2 x^2-x-20 cannot be factored so you would have to use quadratic formula -b+-√b^2-4ac ---------------- 2 1+-√1^2-4(1)(-20)=1+-√1+80=1+-√81=9 1+9/2=5 or 1-9/2=-4 (x+5)(x-4)

OpenStudy (anonymous):

(x+4)(x-5) because if it were the other way it would end up x^2+x-20 not x^2-x-20

OpenStudy (anonymous):

The equation can be factored into 3(x-5)(x+4). First, factor out the common factor 3, do not divide. Since the remaining expression is \[x ^{2}-x-20\] this represents a trinomial with leading coefficient of 1. So, if it is going to factor, you need the factors of a -20 which add up to the coefficient of the middle term. The coefficient of the middle term is a -1. So the factors of -20 which add to a -1 are -5 and 4. Therefore, your remaining polynomial factors as (x-5)(x+4). And thus your entire expression, written in complete factored form is 3(x-5)(x+4).